Could someone please assist me by solving this logarithmic equation step-by-step for future reference. Thank you.
log(7y + 1) = 2log(y+3)-log2
log(7y + 1) = log(y+3)^2-log2
log(7y + 1) = log(y+3)^2/2
you can write
(7y + 1) = (y+3)^2/2
2(7y + 1) = y^2+6y+9
14y+2=y^2+6y+9
y^2-8y+7=0
y1=(8+rootsquare(36))/2
=(8+6)/2
=7
y2=(8-rootsquare(36))/2
=(8-6)/2
=1
so (y1,y2) are (7,1)

Found 2 solutions by scott8148, bimanewton:
Answer by scott8148(6628)

handy rules for logs:
adding logs is multiplying numbers
subtracting logs is dividing numbers
multiplying logs is exponentiation
dividing logs is taking roots

with those in mind ... 7y+1=((y+3)^2)/2 ... 14y+2=y^2+6y+9

y^2-8y+7=0 ... (y-7)(y-1)=0

Answer by bimanewton(8)